(x2 + y2)dx + x(x − 2y)dy = 0


1
Expert's answer
2021-04-29T17:58:47-0400

(x²+ y²)dx + x(x − 2y)dy = 0

=> "\frac{dy}{dx}=\frac{x^2+y^2}{2xy-x²}"

This is a homogeneous differential equation

Let us put y = vx and so "\frac{dy}{dx}=v+ x\frac{dv}{dx}"

The equation transforms to

"v+ x\frac{dv}{dx}" = "\frac{x^2+v^2x^2}{2vx^2-x²}"

=> "v+ x\frac{dv}{dx}" = "\frac{1+v^2}{2v-1}"

=> "x\frac{dv}{dx}" = "\frac{1+v^2}{2v-1} - v"

=> "x\frac{dv}{dx}" = "\frac{1+v-v^2}{2v-1}"

=> "\frac{2v-1} {1+v-v^2} dv" = "\frac{dx} {x}"

=> "\frac{2v-1} {v^2-v-1} dv" = "-\frac{dx} {x}"

=> "\int\frac{2v-1} {v^2-v-1} dv =" "-\int\frac{dx} {x}" +ln|C|

=> ln|v²-v-1} = - ln|x| + ln|C| [since "\int\frac{f'(x)}{f(x)} dx = ln f(x)]"

=> ln|v²-v-1| + ln|x| = ln|C|

=> ln|v²x-vx-x} = ln|C|

=> "ln| \frac{y²}{x} -y -x| = ln|C|"

=> "ln| \frac{y²-xy -x²}{x}| = ln|C|"

=> "| \frac{y²-xy -x²}{x}| = |C|"

=> "(y\u00b2-xy -x\u00b2)\u00b2= C\u00b2x\u00b2"